OpenStudy (anonymous):

Medal A one-serving size bag of potato chips contains 75 calories from fat, which represents 60% of the total calories of the chips in the bag. How many calories are there in this bag of chips? ? calories

OpenStudy (texaschic101):

well...75 calories is 60%....so we need to find 100% 75/60 = x/100 --- (75 cal to 60% = x cal to 100%) cross multiply (60)(x) = (75)(100) 60x = 7500 x = 7500/60 x = 125 so if 75 calories = 60%, then 125 calories = 100%

OpenStudy (anonymous):

so your answer is 125 cal

OpenStudy (texaschic101):

yes it is :)

OpenStudy (anonymous):

thank you sooooo much :)

OpenStudy (texaschic101):

no problem :)